OpenStudy (anonymous):

Integrate (x^2-x+6)/(x^3+3x)

OpenStudy (lgbasallote):

have you tried using rational expressions?

OpenStudy (anonymous):

I believe it's meant to be done by partial fractions but I am having a hard time of it

OpenStudy (anonymous):

x^2-x+6 = a(x)+b(x^2+3) is as far as I have got.

OpenStudy (lgbasallote):

x(x^2+ 3) so partial fraction is: A/x + (Bx +C)/x^2+3 x^2 -x +6 = Ax^2 + 3A + Bx^2 + Cx Gaussian method: coeefficients of x^2: 1 = A + B coefficients of x -1 = C coefficients of constant: 6 = 3A A = 2 B = -1 so your equation is 2/x + (-1 -x)/x+3 so... 2/x - (1+x)/(x+3) can you do it from there?

OpenStudy (anonymous):

Can I ask why is it (Bx+C)? Instead of just B/x^2+3?

OpenStudy (lgbasallote):

because x^2 + 3 is an irreducible quadratic fractions (Case III of partial fractions) If you have irreducible quadratic functions, the numerator is composed of Bx + C or any other distinct letters

OpenStudy (anonymous):

Thanks again, I can see that part in my book but it's not very well explained. Thanks !